OpenStudy (anonymous):

For what value "c" is the function continuous on the interval: (-infinity, infinity)? f(x) = {cx^2 + 8x, when x < 5 f(x) = {x^3 - cx, when x ≥ 5 Assume thats just one big bracket I got c = 8/3. Is this right?

OpenStudy (anonymous):

) An object is thrown upward from the edge of a tall building with a velocity of 10m/s. Where will the object be 3s after it is thrown? Take g = 10m/s^2

OpenStudy (anonymous):

pls help me

OpenStudy (turingtest):

@successboye79 I did, look at your post

OpenStudy (anonymous):

Is this right?

OpenStudy (turingtest):

I'm trying to figure it out...

OpenStudy (anonymous):

set them equal to each other and solve for c

OpenStudy (mrsmarsh):

8/3 is not correct

OpenStudy (anonymous):

What I did was I equated the one-sided limits together

OpenStudy (anonymous):

So limit 5 as x approaches it from the left side = limit 5 as x approaches it from the right side That would be: cx^3 + 8x = x^3 - cx

OpenStudy (anonymous):

c(5)^3 + 8(5) = (5)^3 - c(5)

OpenStudy (anonymous):

25c + 40 = 125 - 5c

OpenStudy (anonymous):

20c = 40 - 125

OpenStudy (anonymous):

20c = 85

OpenStudy (anonymous):

85/20 = c?

OpenStudy (anonymous):

Is this correct?

OpenStudy (turingtest):

\[cx^2+8x=x^3-cx\]\[cx^2+cx=cx(x+1)=x^3-8x=x(x^2-8)\]\[c=(x^2-8)/(x+1)\]we need this to be continuous at x=5 so...\[c=(25-8)/(5+1)=17/6\]

OpenStudy (turingtest):

that took me way too long, I'm just waking up... :/